OpenStudy (anonymous):

Help anyone? determine a function f(x) = mx+b for the graph whose line is parallel to x - y = 12 and passes through the point (5,-1)

OpenStudy (anonymous):

y = -x-12 slope, m = -1 y - y1 = m ( x - x1) y - (-1) = -1(x-5) y + 2 = -1x + 5 y = -1x + 3

OpenStudy (anonymous):

oops, y - (-1) = 1 y = -1x + 4

OpenStudy (anonymous):

Haiz your first part i believe is wrong x-y=12 x=y+12 x-12 = y

OpenStudy (anonymous):

m=1 ... parallel lines have same slope (x1,y1)=(5,-1) y-y1=m(x-x1) y+1=1(x-5) y+1=x-5 y=x-4

OpenStudy (anonymous):

y=x-6

OpenStudy (anonymous):

oh hah ok

OpenStudy (anonymous):

your line is y=x-6 , my bad check (5,-1) -1=5-6=-1 check

OpenStudy (anonymous):

they need an edit button

OpenStudy (anonymous):

i see how its done. thanks a lot!

OpenStudy (anonymous):

the error is easy to see y+1=x-5 y=x-6